# Towards autonomous photogrammetric forest inventory using a lightweight under-canopy robotic drone

### v1

@@ Abstract
Drones are increasingly used in forestry to capture high-resolution remote sensing data. While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting. Inside dense forests, reliance on the Global Navigation Satellite System (GNSS) for localization is not feasible. Additionally, the drone must autonomously adjust its flight path to avoid collisions. Recently, advancements in robotics have enabled autonomous drone flights in GNSS-denied obstacle-rich areas. In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open-source methods and validating its performance for data collection inside forests. The autonomous flight capability was evaluated through multiple test flights in two boreal forest test sites. The tree parameter estimation capability was studied by conducting diameter at breast height (DBH) estimation using onboard stereo camera data and photogrammetric methods. The prototype conducted flights in selected challenging forest environments, and the experiments showed excellent performance in forest reconstruction with a miniaturized stereoscopic photogrammetric system. The stem detection algorithm managed to identify 79.31% of the stems. The DBH estimation had a root mean square error (RMSE) of 3.33 cm (12.79%) and a bias of 1.01 cm (3.87%) across all trees. For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74%), and the bias was 0.13 cm (0.64%). When considering the overall performance in terms of DBH accuracy, autonomy, and forest complexity, the proposed approach was superior compared to methods proposed in the scientific literature. Results provided valuable insights into autonomous forest reconstruction using drones, and several further development topics were proposed.
@@
articletype: ARTICLE PREPRINT
@@ 1 Introduction
Over the last decade, the use of drones in forestry has rapidly increased, offering a practical solution for forest data collection both above and inside the forest canopy (Puliti et al. 2015; Hyyppä et al. 2020a; Liang et al. 2019). Modern drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems rely exclusively on Global Navigation Satellite Systems (GNSS) signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al. 2010). Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path. Concequently, under-canopy data collection with drones is still primarily controlled manually by human pilots (e.g. Hyyppä et al. 2020a; Liang et al. 2019; Krisanski, Taskhiri, and Turner 2020).
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments. In recent years, several solutions utilizing various sensor setups and algorithms for autonomous navigation under forest canopies have been proposed (e.g. Liu et al. 2022; Loquercio et al. 2021; Zhou et al. 2022). However, these methods are typically validated by reporting a single test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests (Loquercio et al. 2021; Ren et al. 2022; Liu, Ren, and Zhang 2024). Therefore, a significant research gap remains regarding the applicability and reliablity of these algorithms in diverse forest environments.
The autonomous under-canopy drone solutions proposed in the literature typically focus on aerial robotics and primarily utilize the collected sensor data for navigation. However, recent research has begun to explore boarder applications. For example, Liu et al. (2022) employed a semantic simultaneous localization and mapping (SLAM) algorithm to calculate the number of trees in the explored area. Similarly, Liang et al. (2024) demonstrated the use of autonomous drones for in situ forest observations under the canopy. While their work estimated tree attributes based on data collected during two test flights, it did not evaluate the reliability of autonomous flight itself. Consequently, a research gap remains regarding the borader application of autonomous drones for forestry applications.
@@
Remote sensing measurements within forests are increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al. 2007) and mobile solutions e.g. handheld devices (Ryding et al. 2015; Bauwens et al. 2016; Marselis et al. 2016; Balenović et al. 2021), backpacks (Liang et al. 2014b, 2018; Hyyppä et al. 2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al. 2018; Bienert et al. 2018) or manually operated drones (Chisholm et al. 2013; Brede et al. 2017; Wieser et al. 2017; Hyyppä et al. 2020a). Similarly, the use of structure-from-motion (SFM) based point clouds for forest mapping has been demonstrated using different camera solutions (Liang et al. 2014a; Wallace et al. 2016; Kuželka and Surovỳ 2018; Piermattei et al. 2019). Combining these techniques with autonomously navigating drones inside forests represents a highly relevant research opportunity.
To address these gaps and opportunities, this study introduces a novel approach for autonomous data collection and object characterization under unstructured forest canopies. The primary objectives were to assess the performance of an autonomously flying drone in a boreal forest and to demonstrate its capability in measuring tree characteristics. A prototype autonomously flying drone was developed using state-of-the-art open-source algorithms, followed by multiple flight tests to evaluate its performance and suitability for boreal forests. Image data collected by the low-cost onboard stereo-camera was stored for further analysis. The dataset was post-processed using SfM photogrammetry to improve the georeferencing and to generate dense 3D point clouds. Finally, tree detection and diameter at breast height (DBH) estimation were performed on the point clouds, assessing the potential of the low-cost autonomous system for forest parameter estimation.
@@
The study is structured as follows. Section 2 reviews recent open-source solutions for autonomous drone flights inside forests and state-of-the-art methods forest parameter estimation using drones flying inside forests. Section 3 describes the algorithms, hardware, experiments, and methods for autonomous navigation, tree detection, and DBH estimation. Experimental results are presented in Section 4. Section 5 discusses the results, potential improvements, and concludes the study.
@@ 2 Related work
2.1 Autonomous flying inside forests
Autonomous navigation in obstacle-rich environments has been an actively researched topic in robotics during the last decades. Whereas the majority of the research has tried to tackle the 2D path planning problem for autonomous ground vehicles, in recent years increasing number of methods for autonomous path planning and obstacle avoidance have also been published for drones. Although many of the methods proposed in the literature are only tested in simulations or indoor conditions, some of the methods have been also successfully flight-tested inside forests. Some of the authors have also open-sourced their methods to encourage the development of the field. In this section, recent open-source methods successfully flight-tested inside a forest are summarized.
Campos-Macías et al. (2021) proposed a solution that focuses on computational efficiency that can be run in real time with a low-power CPU. The solution used a stereo tracking camera to run Visual-Inertial Odometry (VIO) for pose estimation and a point cloud from an active stereo depth camera to create a 3D voxel map. Collision-free trajectories were planned to the free regions on the map. In addition to multiple indoor environments, the system was tested in a sparse mixed forest. However, the flight distance in the forest test was only 11.58 m.
Liu et al. (2022) proposed a method that aims at large-scale autonomous flights in forest environments. The proposed system used an inertial measurement unit (IMU) and a stereo camera to run VIO and 3D light detection and ranging (LiDAR) to run a semantic simultaneous localization and mapping (SLAM) algorithm. Semantic SLAM was developed to correct drift in the VIO estimate. The system was reported to fly an 800 m long autonomous flight, but the semantic SLAM algorithm was only tested in manual flight. The documented flight tests were performed in a North American pine forest. The forest was relatively dense, but it had a minimal amount of low branches and understory vegetation.
Ren et al. (2022) proposed another LiDAR-based method that generates sphere-shaped safe flight corridors from the point clouds. The method was targeting for high-speed flying in unknown environments. FAST-LIO2 (Xu et al. 2022) was utilized in the pose estimation. In total, 12 test flights with varying flight speeds were performed in a sparse Chinese deciduous forest. The maximum flight speed was 13.7 m/s and 11 of 12 test flights were successful, but the trajectory lengths were not reported. Liu, Ren, and Zhang (2024) have also proposed another method for the same LiDAR-based platform targeting especially safe navigation in the presence of dynamic obstacles and external disturbances with an integrated planning and control framework. Planning was performed within a local grid map and FAST-LIO2 (Xu et al. 2022) was utilized in the pose estimation. In addition to indoor environments, the system was also tested with 10 successful flights at varying speeds in the same forest as (Ren et al. 2022). The maximum flight speed was 5.86 m/s, and the reported trajectory length was 58.64 m.
Instead of using a single drone, the solution proposed by Ahmad et al. (2022) focused on the navigation of drone swarms. The proposed approach is fully decentralized and tries to mimic the collective navigation behavior of natural flocks and swarms. The physical drone hardware used in the tests incorporated a LiDAR for running the SLAM algorithm by Kohlbrecher et al. (2011), a laser rangefinder as an altimeter, and an Ultra Violet Direction And Ranging (UVDAR) system for relative localization between the drones of the swarm. The approach was tested in a swarm flight inside a forest, but the test forest was very clean containing no low branches or understory vegetation. The flight distance was approximately 40 m.
Zhou et al. (2022) also proposed a solution concentrating on swarm flights. In the proposed solution the drones use VIO for localization and probabilistic maps for obstacle avoidance. Swarm members also share their trajectories over a wireless network and correct the VIO drift based on the relative distances between the drones measured with Ultra Wide Band (UWB) sensors. The proposed solution was tested with palm-sized drones which used only one stereo camera to run both mapping and VIO. In the flight tests, a swarm of drones was successfully flying approximately 80 m through a dense bamboo forest.
The solution proposed by Loquercio et al. (2021) replaced different sensing, mapping, and planning subtasks with a single neural network directly mapping depth images and a pose estimate from VIO to a feasible collision-free trajectory for the drone. The system was trained entirely with simulated data, and the scope of the research was high-speed autonomous flying. Real-world flight tests were performed with a drone using an active stereo depth camera to produce the depth images, and a stereo tracking camera to run VIO. Flight tests were performed in various environments containing also Central-European mixed forests. The forest densities were not reported, but the forests were relatively sparse with a low amount of understory vegetation. The evaluated trajectories were a 40 m long straight line and a circle with a 6 m radius, and multiple test flights were conducted with flight speeds varying from 3 to 10 m/s. The flight success rate was 100 % with a flight speed of 3 m/s and 5 m/s, 80 % with a flight speed of 7 m/s, and 60 % with a flight speed of 10 m/s
Nguyen et al. (2023) proposed also a deep-learning-based method that utilizes neural networks to directly map depth images and partial pose estimates of the drone to a series of actions leading to a collision-free trajectory. The proposed system was also trained entirely in a simulator. In addition to the indoor flight tests, the performance assessment was reported for a single 60 m long trajectory through a dense Finnish pinewood forest making the test environment the most similar to the target environments in this study.
@@ 2.2 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important tree attributes to be estimated since it provides a measure for the size of a tree and also enables the prediction of the total stem volume (Laasasenaho 1982). During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al. (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al. (2018); Piermattei et al. (2019); Xu, Shen, and Cao (2023)) and above-canopy photogrammetric systems (e.g., Wallace et al. (2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023). If the operation of such under-canopy drone systems can be automated in the future, they would provide a truly efficient solution for the acquisition of accurate forest data.
Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method. The studied plots were sparse with only 270-290 trees/ha. Krisanski, Taskhiri, and Turner (2020) performed manual flights of a drone system inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with a RMSE of 1.1-2.1 cm. Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small forest plots using SfM point clouds acquired with an under-canopy flying drone system. The drone system was able to detect obstacles in the horizontal direction, but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction. They also demonstrated stem volume estimates of individual trees with a RMSE of 15.8-17.2%.
@@
In addition to photogrammetric approaches, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al. 2015; Bauwens et al. 2016; Cabo et al. 2018; Chen et al. 2019; Hyyppä et al. 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation. Even though a LiDAR can be mounted to a variety of other mobile platforms besides a drone, the focus is on the use of under-canopy drones for DBH estimation in this brief literature review.
@@
Chisholm et al. (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually-piloted drone flying under the forest canopy. They detected 12 trees in the vicinity of the drone trajectory and reported a RMSE of 25.1% for the DBH estimates of the detected trees.
@@
Hyyppä et al. (2020a) collected high-quality point cloud data using a Kaarta Stencil-1 laser scanner mounted horizontally at the bottom of a manually-piloted drone. They demonstrated DBH estimates with a RMSE of 0.6-0.9 cm (2%-3%) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. As the laser scanner was mounted horizontally in their study, the limited field of view made it difficult to obtain direct measurements of tree height. In a follow-up study by Hyyppä et al. (2021), this issue was resolved by instead integrating a rotating laser scanner to an under-canopy flying drone, which resulted in accurate measurements for both stem diameter and tree height. Wang et al. (2021) combined data from above- and under-canopy flights of a manually-piloted drone, resulting in an RMSE of 7.95 cm (33.35%) and a bias of 3.72 cm (15.61%) for the DBH.
@@
Muhojoki et al. (2024) compared the accuracy of the estimated tree attributes from several LiDAR-based mobile systems, including two under-canopy drone systems, a system by Hovermap (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a system by Deep Forestry (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). Both drones were flown manually in Muhojoki et al. (2024) to collect data in sparse density forest plots. The RMSE and bias of DBH estimates for the Deep Forestry’s and Hovermap’s drones, with bias removal method introduced by Muhojoki et al. (2024), averaged RMSE of 1.5 cm (6.7 %) and bias of -0.7 cm (3.0 %) and RMSE of 1.6 cm (6.9 %) and bias of -0.1 cm (-0.6 %), respectively. Without using bias removal method, RMSE and bias for both drones were significantly higher. Muhojoki et al. (2024) also demonstrated that the under-canopy drone systems can provide roughly as accurate tree attributes as other under-canopy MLS methods, such as handheld LiDARs.
@@
Recently, Liang et al. (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable the autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environment conditions. When the flight path covered most of the test plot, they achieved an RMSE of 5.1 cm for the DBH of individual trees, and a similar RMSE for the stem curve.
@@
In addition to under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al. 2017; Brede et al. 2017; Kuželka, Slavík, and Surovỳ 2020; Vandendaele et al. 2021; Feng et al. 2022).
A summary of relevant studies on under canopy forest inventory studies is presented in Table 1.

### v2

@@ Abstract
Drones are increasingly used in forestry to capture high-resolution remote sensing data, supporting enhanced monitoring, assessment, and decision-making processes. While operations above the forest canopy are already highly automated, flying inside forests remains challenging, primarily relying on manual piloting. Inside dense forests, reliance on the Global Navigation Satellite System (GNSS) for localization is not feasible. Additionally, the drone must autonomously adjust its flight path to avoid collisions. Recently, advancements in robotics have enabled autonomous drone flights in GNSS-denied obstacle-rich areas. In this article, a step towards autonomous forest data collection is taken by building a prototype of a robotic under-canopy drone utilizing state-of-the-art open-source methods and validating its performance for data collection inside forests. Specifically, the study focused on camera-based autonomous flight under the forest canopy and photogrammetric post-processing of the data collected with the low-cost onboard stereo camera. The autonomous flight capability of the prototype was evaluated through multiple test flights at boreal forests. The tree parameter estimation capability was studied by performing diameter at breast height (DBH) estimation. The prototype successfully carried out flights in selected challenging forest environments, and the experiments showed excellent performance in forest 3D modeling with a miniaturized stereoscopic photogrammetric system. The DBH estimation achieved a root mean square error (RMSE) of 3.33 cm (12.79 %) across all trees. For trees with a DBH less than 30 cm, the RMSE was 1.16 cm (5.74 %). The results provide valuable insights into autonomous under-canopy forest mapping and highlight the critical next steps for advancing lightweight robotic drone systems for mapping complex forest environments.
@@
articletype: ARTICLE PREPRINT - Word Count 10048
@@ 1 Introduction
Over the last decades, remote sensing measurements within forests have been increasingly used to capture detailed forest information. Terrestrial laser scanning (TLS) made a breakthrough in the early 2000s, followed by the development of various solutions, including static laser scanning using tripods (Bienert et al. 2007) and mobile solutions, e.g., handheld devices (Ryding et al. 2015; Bauwens et al. 2016; Marselis et al. 2016; Balenović et al. 2021), backpacks (Liang et al. 2014b, 2018; Hyyppä et al. 2020b), ground-based vehicles (Forsman, Holmgren, and Olofsson 2016; Liang et al. 2018; Bienert et al. 2018) or drones (Chisholm et al. 2013; Brede et al. 2017; Wieser et al. 2017; Hyyppä et al. 2020a). Similarly, the use of structure-from-motion (SFM) based point clouds for forest mapping has been demonstrated using different camera solutions (e.g., Liang et al. 2014a; Wallace et al. 2016; Kuželka and Surovỳ 2018; Piermattei et al. 2019). Among these methods, drone-based solutions offer several benefits, including the ability to move in areas with challenging terrain and the high potential for automatization.
Hence, the use of drones in forestry research has rapidly increased, and they offer a practical solution for forest data collection above the forest canopy (Puliti et al. 2015; Tuominen et al. 2017). Modern commercial drones are highly automated and can execute data collection missions highly autonomously in open-air environments using autopilot systems. However, their positioning systems typically rely exclusively on GNSS signals, limiting autonomous flights to open areas where these signals are reliable. In dense forests, GNSS localization becomes unreliable due to signal blockages and multipath effects caused by trees (Schubert et al. 2010). Additionally, navigating inside forests requires the capability to autonomously avoid obstacles such as trees and bushes along the flight path. Consequently, drone data collection flights under the forest canopy are still primarily controlled manually by human pilots (e.g., Hyyppä et al. 2020a; Liang et al. 2019; Krisanski, Taskhiri, and Turner 2020; Prabhu et al. 2024).
Recent advances in robotics and the enhanced computational capabilities of small embedded computers have enabled drone autonomy even in GNSS-denied, cluttered environments. Although technology is still in the early stage of development, some recent studies have proposed various methods that have been successfully tested in forest environments (e.g., Liu et al. 2022; Loquercio et al. 2021; Zhou et al. 2022). However, the proposed methods are typically validated by reporting a single successful test flight in a specific forest environment. Most studies assess the effects of varying object density only in simulators, with only a couple of the methods tested through multiple test flights inside real forests.
Moreover, most of the studies proposing methods for autonomous under-canopy drone trajectory planning have the scope solely on aerial robotics, and the sensor data is primarily utilized only for navigation. Only a few pioneering studies have yet explored the potential of custom robotic under-canopy drone systems in forestry applications (Liu et al. 2022; Liang et al. 2024; Cheng et al. 2024). Notably, all of these studies have utilized Ouster OS1 3D LiDAR (Ouster, San Francisco, CA, USA) as the main sensor. Whereas LiDARs have their strengths, such as 360° horizontal field of view (FOV), long sensing range, and robustness against varying illumination, they tend to be heavier and have higher power consumption than cameras (Reddy Cenkeramaddi et al. 2020). In the case of drone development, where optimizing the weight and battery consumption play a key role, these factors lead to bigger and heavier drone platforms and flight batteries requiring more space between the obstacles to fly safely inside the forest. Moreover, LiDARs with suitable quality are more expensive than camera systems with built-in 3D imaging capability, leading to increased building cost. The data collected with the LiDARs is also limited to point clouds only, whereas the images collected with cameras can also be used for image-based analysis.
Hence, although some LiDAR-based robotic drone systems have been successfully applied to under-canopy mapping of forest parameters, the potential of smaller and cheaper camera-based systems remains largely unexplored. Furthermore, the robotic drone systems proposed in the literature are typically not evaluated through multiple flight tests in different forest environments leaving the reliability and suitability to diverse forest environments as an open question.
Therefore, in this study, a prototype of a small camera-based robotic drone system that can fly autonomously in challenging GNSS-denied forests was built and developed using custom hardware and state-of-the-art robotic methods. The study aimed to tackle the research gaps in flight reliability of camera-based robotic drone systems in various boreal forest environments, as well as the potential of photogrammetric DBH estimation with autonomous drones in such environments. The performance and the reliability of autonomous flight capabilities were evaluated by performing multiple test flights in boreal forests with varying difficulty. In addition, the image data collected with the low-cost onboard stereo camera were processed into point clouds, and the DBH values were estimated for the trees detected from the dataset. To the best of our knowledge, this study is the first to empirically validate a miniaturized, autonomous, camera-based robotic drone for under-canopy forest reconstruction.
Specifically, the study focused on three research questions. What level of autonomous flight reliability can be achieved in challenging boreal forests with current methods for camera-based robotic drone navigation? What level of DBH estimation accuracy can be achieved with the low-cost onboard stereo-camera data recorded during the flight? What are the remaining bottlenecks to be tackled in follow-up studies in order to achieve large-scale autonomous photogrammetric forest data collecting with miniaturized under-canopy robotic drone systems?
The main contributions of this study can be highlighted as follows:
  • Build and development of a lightweight robotic drone prototype capable of flying within GNSS-denied forest environments, using custom hardware and state-of-the-art robotic algorithms. The reliability of the system was comprehensively evaluated through multiple test flights in boreal forests with varying tree densities and difficulty levels.
  • Combining autonomous flight capabilities to under-canopy photogrammetric forest reconstruction by processing the data from a low-cost onboard stereo-camera to 3D point clouds.
  • Validation of autonomous flight data collection based on stereo cameras for the estimation of forest parameters through DBH estimation, including an evaluation of the impact of various factors on the accuracy of the estimate.
  • Comprehensive discussion of current limitations to large-scale, fully autonomous photogrammetric under-canopy forest monitoring, along with proposed solutions and further research directions.
@@
The rest of this study is structured as follows. Section 2 reviews the latest studies that estimate DBH using drones flying inside forests and recent open-source solutions for autonomous drone flights inside forests. Section 3 describes the algorithms, hardware, experiments, and methods utilized in this study. Experimental results are presented in Section 4. Section 5 discusses the results, and potential improvements, and concludes the study.
@@ 2.1 DBH estimation from under-canopy drone data
In a forest inventory, DBH is one of the most important attributes to estimate, as it provides a measure of the size of a tree and enables the prediction of the total stem volume either alone or combined with the three height (Laasasenaho 1982). Although some of the relevant studies have also estimated tree volume by flying the flight path multiple times with different altitudes, with rotating sensors, or with Lidars with high vertical FOV, only DBH estimation accuracies and methods are included in this brief literature review.
@@
In the existing literature, LiDAR-based methods have been studied extensively for the purpose of efficiently and accurately estimating various tree attributes, such as DBH, tree height, and stem volume. To date, several LiDAR-based mobile platforms have been shown to enable DBH estimation of individual trees with an RMSE of 1-2 cm (Ryding et al. 2015; Bauwens et al. 2016; Cabo et al. 2018; Chen et al. 2019; Hyyppä et al. 2020b, a, 2022), though most studies have focused on relatively sparse forests with little understory vegetation.
@@
Chisholm et al. (2013) were the first to demonstrate DBH measurements from point cloud data collected with a manually piloted drone flying under the forest canopy. The utilized custom-built drone system included Hokuyo UTM-30LX 2D LiDAR (Hokuyo, Osaka, Japan). They detected 12 trees in the vicinity of the drone trajectory and reported an RMSE of 25.1 % for the DBH estimates of the detected trees.
@@
Hyyppä et al. (2020a) collected high-quality point cloud data using a Kaarta Stencil-1 (Kaarta, Pittsburgh, Pennsylvania, USA) laser scanner mounted horizontally at the bottom of a manually piloted drone. They demonstrated DBH estimates with an RMSE of 0.69-0.92 cm (2.2 %-3.1 %) and a bias of 0.29-0.31 cm (1.0 %-1.1 %) on two plots of varying complexity in a boreal forest with a tree density of 410-420 trees/ha. Importantly, their stem diameter estimation algorithm utilized the temporal order of data points to detect stem arcs, thus mitigating distortions caused by slow temporal drifts of the drone trajectory. In a follow-up study, Hyyppä et al. (2020c) compared various mobile laser-scanning methods, including backpack, handheld, Under-Canopy UAV, and Above-Canopy UAV, for forest data collection. The same manually piloted under-canopy drone system was utilized in this study in the same forest plots as in Hyyppä et al. (2020a). In this study, the manually piloted under-canopy drone system yielded an RMSE of 0.6-1.1 cm (2.3 %-3.5 %) and a bias of 0.34-0.12 cm (1.3 %-0.4 %). Wang et al. (2021) combined data from above- and under-canopy flights of a manually-piloted drone, resulting in an RMSE of 7.95 cm (33.35 %) and a bias of 3.72 cm (15.61 %) for the DBH.
@@
Muhojoki et al. (2024) compared the accuracy of the estimated tree attributes from several LiDAR-based mobile systems, including two under-canopy drone systems, a system by Hovermap (2020 version, Emesent, Milton, Australia) equipped with a Velodyne VLP-16 laser scanner (Velodyne Lidar, San Jose, CA, USA) and a system by Deep Forestry (Deep Forestry, Uppsala, Sweden) utilizing an Ouster OS0-32 Rev. 5 laser scanner (Ouster, San Francisco, CA, USA). Deep Forestry drone was flown manually, but the Hovermap drone system was able to fly autonomously in sparse forests where the distance between the trees was at least 3 m. Both drone systems were used to collect point clouds under the canopy of three sparse forest plots. The RMSE and bias of DBH estimates for Deep Forestry and Hovermap drones, with bias removal method introduced by Muhojoki et al. (2024), averaged RMSE of 1.5 cm (6.7 %) and bias of -0.7 cm (3.0 %) and RMSE of 1.6 cm (6.9 %) and bias of -0.1 cm (-0.6 %), respectively. Without using the bias removal method, RMSE and bias for both drones were significantly higher.
@@
In addition to the manually controlled drones and large autonomous commercial drone systems, recently a few pioneering studies have applied smaller custom robotic drone platforms with a high automation level for LiDAR-based under-canopy DBH estimation. Liang et al. (2024) applied an autonomously operated under-canopy drone system to collect laser scanning data inside a 50 m $\times$ 30 m forest plot with a tree density of approximately 1000 trees/ha. To enable autonomous flight, a predefined global path was planned, to which the drone system performed real-time adjustments during the flight according to real-time environmental conditions. When the flight path covered most of the test plot, they achieved an RMSE of 5.13 cm (22.01 %) and bias of 3.40 cm (14.62 %) for the DBH of individual trees.
Liu et al. (2022) employed a LiDAR-based semantic lidar odometry and mapping framework (SLOAM) (Chen et al. 2020) algorithm to calculate the number of trees in the explored area. Although in this study the autonomous flight was demonstrated without the SLOAM algorithm and the tree number calculation was only demonstrated with one manually-piloted test flight, the same drone system has later been used in various follow-up studies for tree parameter estimation. Prabhu et al. (2024) extended the system to have 3D trajectory planning capability and used the platform to demonstrate metric-semantic map merging between multiple robotic drones as well as DBH and tree trunk diameter profile estimation with the onboard LiDAR data. The DBH values were estimated for 20 trees located in two different test plots with RMSE of 1.67 cm (7.42 %). However, the DBH data-collecting flights were performed with manual flight mode. Cheng et al. (2024) used the same drone platform to collect a dataset for LiDAR-based mapping of trees. In addition to data from orchards, the dataset contains the onboard LiDAR data from two forest areas. Although the drone was manually piloted in one of these forests, the data collection at the other forest consisting of loblolly pine trees was performed in autonomous flight mode. They also estimated the DBH value for 77 trees from the LiDAR data collected during the autonomous flights achieving an RMSE of 3.8 cm (12.3 %) by offline afterprocessing and 9.8 cm (31.8 %) by online estimation with SLOAM. With manually piloted data collection in the second test forest consisting of 20 trees with varying species, the RMSEs were 2.0 cm (7.5 %) with offline processing and 7.5 cm (27.5 %) with online processing.
@@
Besides the under-canopy flying drones, LiDARs have also been integrated to above-canopy flying drones capable of direct stem diameter measurements, though typically with a significantly higher RMSE and a lower tree detection rate due to the larger measurement distance and the upper canopy obstructing the tree stems (Wieser et al. 2017; Brede et al. 2017; Kuželka, Slavík, and Surovỳ 2020; Vandendaele et al. 2021; Feng et al. 2022).
During the past decade, photogrammetric systems producing SfM point clouds have been suggested to provide an efficient, low-cost alternative to traditional field measurements of tree attributes, such as DBH. Even though several studies have investigated terrestrial photogrammetric systems (e.g., Liang et al. (2014a); Surovỳ, Yoshimoto, and Panagiotidis (2016); Mokroš et al. (2018); Piermattei et al. (2019); Xu, Shen, and Cao (2023); Gao and Kan (2022)) and above-canopy photogrammetric systems (e.g., Wallace et al. (2016); Ye, van Leeuwen, and Nyktas (2019)) for tree attribute estimation, only relatively few studies have tested the use of under-canopy drone systems for the collection of SfM point clouds enabling tree attribute measurements (Kuželka and Surovỳ 2018; Krisanski, Taskhiri, and Turner 2020; Shimabuku, Konoshima, and Ota 2023; He et al. 2025).
Kuželka and Surovỳ (2018) manually operated a drone system based on a DJI Phantom 4 Pro (DJI, Shenzhen, China) within two forest plots to produce low-noise photogrammetric point clouds, from which the DBH of individual trees could be estimated with an RMSE between 2.6 cm and 5.3 cm depending on the plot and the applied analysis method. The studied plots were sparse with only 270-290 trees/ha. Krisanski, Taskhiri, and Turner (2020) performed manual flights with a DJI Phantom 4 Pro inside two forest plots with a stem density of approximately 1000 trees/ha to acquire high-density SfM point cloud data, from which the DBH of individual trees could be measured with an RMSE of 1.1-2.1 cm.
Shimabuku, Konoshima, and Ota (2023) achieved an RMSE of 0.4-0.7 cm for DBH estimation on two small subtropical forest plots using SfM point clouds acquired with an under-canopy flying DJI Mini 3 Pro drone (DJI, Shenzhen, China). The first plot containing 18 trees was clean containing a minimal amount of low-hanging branches and understory vegetation. The second test plot was denser and closer to a natural subtropical forest environment but contained only 7 trees. The drone system was able to detect obstacles in the horizontal direction but needed manual piloting to avoid obstacles in the vertical direction and to guide the drone back to the intended flight path after automatic obstacle avoidance in the horizontal direction.
Recently, He et al. (2025) mounted an additional 360° FOV Insta360 camera (Arashi Vision, Shenzhen, China) to the same LiDAR-based autonomous drone platform used in Liu et al. (2022) and post-processed the photogrammetric data to 3D point clouds. Whether the flights were performed in autonomous flight mode or with manually piloted mode was not mentioned in the study. Their study evaluated DBH for 43 trees with varying tree species and the reported median relative DBH error was 11.5 % estimated from the photogrammetric point clouds.
A summary of the results obtained in the relevant studies on under-canopy forest inventory is presented in Table 1. In summary of the related work, most of the studies for under-canopy drone-based DBH estimation have been performed with manually piloted drones. Although a few studies have explored the potential of custom robotic drone systems capable of flying without commands from the pilot during the flight, all of these systems have been based on LiDARs. All of the studies based on photogrammetry and SfM point clouds have utilized commercial drone systems, which have required manual piloting during the data collection.